1. Assume the number of employed people in Pakistan is 22.5 million in 2020. The number of unemployed people during 2020 in Pakistan is 6 million, and the labor force participation rate is 60 percent.                                                            (1.5 Marks)

 

i.                   What is unemployment rate, size of the labor force and working age-population in the year of 2020 in Pakistan?


  1. In each of the following cases, does the increase in prices affect CPI, GDP deflator, or both in Pakistan? Analyze and justify your answer.                                           (2 Marks)

 

i.    The price of imported capital goods used in the production of sugar has increased.

ii.  The price of Tapal tea has increased.

 

  1. Analyze and explain the effect of the minimum wage law on the unemployment rate. Explain your answer by using graph.                                                             (1.5 Marks)
1
Expert's answer
2021-01-28T06:54:27-0500

Solution:

1.). i). Unemployment rate = "\frac{Total \;Number \;of unemployed\; people}{The\; total\; labor \;force} \times 100"

Total labor force = The number of employed + the number of unemployed

= 22.5 + 6 = 28.5 Million

Unemployment rate = "\frac{6}{28.5} \times 100= 21.1\%"


Size of the labor force = Total number of employed people + Total number of unemployed people

= 22.5 + 6 = 31.5 Million


Working age population = Total number of people eligible to work

Labor force participation rate ="\frac{labor force}{working \;age \;population}"


60% = "\frac{31.5}{working \;age \;population}"


Working age population = "\frac{31.5}{60\%} = 52.5 \;Million"

 

2.). i.). The increase in the price of imported capital goods will only affect CPI. This is because the goods imported are included in the CPI basket which is consumed by typical consumers within the country.

 

ii). The increase in Tapal tea price will affect both the CPI and GDP deflator. This is because it makes up a larger part of consumer budgets. The GDP deflator uses as a basket all final goods and services produced in the domestic economy, while the CPI basket also includes goods and services purchased by normal consumers. Therefore, changes in the price of domestically produced consumer goods affect the GDP deflator and CPI more.

 

3.). Raising the minimum wage has a negative impact on unemployment levels. Increasing the minimum wage can lead to increased unemployment, depending on the wage increase, as employers would seek automation as opposed to hiring workers.


This is depicted in the below graph: